A bus covers a certain distance in 3hours 2minutes at an average speed of 45km per hrs.how much time will it take to cover the same distance at speed of 50 km per hours?
Answer by ankor@dixie-net.com(22740)

A bus covers a certain distance in 3hours 2minutes at an average speed of 45km per hrs.
Change 3 hrs 2 min to hrs. 3 + 2/60 = 3.03 hrs
Find how far the bus went in this time at 45 km/hr
d = 3.03 * 45 = 136.5 km
:
how much time will it take to cover the same distance at speed of 50 km per hours?
Time = dist/speed
t = 136.5/50
t = 2.73 hrs which is 2 + .73(60) = 2 hrs 44 min